status reports 19-September-2014 07:48:35

Herschel Status Report - September 2010

27 September 2010

Report for period 24 August to 22 September 2010

Mission operations of the Herschel space observatory continued nominally during the reporting period, with the spacecraft and subsystems all performing as expected.

Spacecraft

The second Direct Liquid Content Measurement will be performed in early November 2010, approximately one year after the first such measurement on 25 November 2009. This measurement assesses the remaining amount of liquid helium (LHe) for the spacecraft's active cooling system, which directly determines Herschel's lifetime for scientific observations. Although the measurement will not establish the exact date when Herschel will run out of liquid helium the measurement results will confirm qualitatively the current situation as compared to the predicted LHe consumption.

Star tracker warm pixels
During the last quarter of 2010 the "bad pixel" table identifying "warm" pixels in the star tracker CCD is expected to fill up to 100% of its capacity (128 pixels). At that time the algorithm identifying such pixels will be adapted to cope with the new situation.

Measures to further lower the CCD temperature are in place to limit the occurrence of "warm" pixels, should this be necessary (see the May 2010 status report for the introduction of these measures). They will only be invoked if a significant increase is seen in the number of artefacts ("speed bumps") in mapping observations caused by "warm" pixels.

Payload

Except for the occasional single event upsets (SEUs), all three instruments continue to operate nominally in the Routine Science Phase.

A few single event upset (SEU) corruptions of SPIRE Instrument Control Unit memory locations have been noted (similar to those for the Instrument Control Unit of HIFI). All but one of those seen so far have been in unused memory or memory not affecting instrument performance. On-board and on-ground operational measures and scheduling procedures have been put in place to capture these and to address, through appropriate planning responses, the very small risk that an SEU might threaten instrument safety.

PACS photometry operations are nominal, and so are the SPIRE/PACS parallel operations.

The PACS anomaly of spontaneous changes in some reference voltage settings (see the July/August status report) is thought to be caused by an instrument hardware issue. The anomaly only affects observations in PACS spectroscopy mode. The frequency with which the disturbances impact the spectroscopy observations could be significantly reduced by updates in the PACS on-board management. Validation of these updates against Flight Spare hardware and their on-board implementation is expected to be completed within a few weeks, during which PACS will only operate in the photometry mode.

No significant change has been observed in the Single Event Upset susceptibility of the HIFI Local Oscillator Control Unit. With 9 months of statistics, on average 2 to 3 events are observed per month. Recovery from such events has become very much routine and operationally the situation is fully under control.

Ground Segment

The technical evaluation of the 576 proposals, received in response to the first in-flight Announcement of Opportunity (AO) for Open Time (OT1) observations, has been concluded on 20 September as planned. More than 40 members of the Time Allocation Committee will meet in the week of Monday 11 October at ESAC, to prepare their recommendation to ESA's Director of Science and Robotic Exploration. The approved OT1 proposals are planned to be announced on 1 November.

Mission Operations
During the reporting period, mission operations have been conducted with the support of the ESA New Norcia ground station. Observational data stored on-board Herschel was received on ground during communication passes lasting approximately three hours.

Archiving
The ground segment is operating nominally. Data products are generated routinely and ingested into the Herschel Science Archive (HSA).

Future Milestones

  • 1 November 2010: Announcement of approved OT1 proposals
  • Early November 2010: Second Direct Liquid Content Measurement


---
Legal disclaimer
This report is based on the Herschel mission manager's report dated 22 September 2010. Please see the copyright section of the legal disclaimer (linked from the home page http://sci.esa.int) for terms of use.


Last Update: 04 January 2011

For further information please contact: SciTech.editorial@esa.int

Status Reports Archive

Related Links

See Also